# https://tet.uni-wuppertal.de

### v1

@@ 1 Introduction
Electromagnetic field problems, where objects are small in comparison to the shortest wave length of the problem, belong to the regime of quasistatic field models, where electromagnetic wave propagation effects are neglected. In addition to the established magneto- and electro-quasistatic field models, electromagneto-quasistatic field models, or so called Darwin-type field models w.r.t. to the original work in [2], have become of interest as quasistatic approximations of the full set of Maxwell equations. These field models are of relevance for applications such as e.g. high-frequency coils or middle frequency transformers (MFTs), where resistive, capacitive and inductive field effects are considered simultaneously. The EMQS models of Darwin-type commonly feature the Darwin-Ampère equation formulated in terms of the magnetic vector potential $\bm{A}$ and the electric scalar potential $\varphi$ as
$\mathrm{curl}(\nu\mathrm{curl}\bm{A})+\kappa\frac{\partial}{\partial t}\bm{A}+\kappa\mathrm{grad}\varphi+\varepsilon\mathrm{grad}\frac{\partial}{\partial t}\varphi=\bm{J}_{\mathrm{S}}.$ (1)
@@
Other EQMS formulations combine (1) with the Maxwell continuity equation [9, 4] which yields symmetric formulations. Combined with the finite integration technique (FIT) [11] or related mimetic discretization schemes this formulation implicitly enforces a Coulomb-type gauge $\nabla\cdot(\varepsilon\frac{\partial}{\partial t}\bm{A})=0.$ Related EMQS formulations as in [12] additionally and explicitly enforce this Coulomb-type gauge within Lagrange multiplier formulations.
A number of EMQS formulations use domain dependent gauge equations and split scalar potentials to enforce a low-frequency stabilization [1, 7].
Port-Hamiltonian systems models [6] are used to maintain physical properties such as energy conservation and/or dissipative inequalities. We consider port-Hamiltonian partial differential-algebraic equation (pH-PDAE) system of the form
${\mathbf{E}}\,\frac{d}{dt}x=({\mathbf{J}}-{\mathbf{R}})\,{\mathbf{z}}(x)+{\mathbf{B}}\,{\mathbf{u}}(t),\quad{\mathbf{y}}(t)={\mathbf{B}}^{\star}\mathbf{z}(x)$ (3)
with a self-adjoint and nonnegative bounded linear operator ${\mathbf{E}}$, a bounded operator ${\mathbf{B}}$, a formally skew-adjoint differential operator ${\mathbf{J}}$ and a dissipative operator ${\mathbf{R}}$. The function $\mathbf{z}$ depends on the inner state variable $\mathbf{x}$ and the input is given by $\mathbf{u}(t)$. If the compatibility relation
@@
holds for the Hamiltonian $H=H(x)$, which represents an energy expression, the dissipativity inequation $\frac{d}{dt}H\leq{\mathbf{y}}^{\top}{\mathbf{u}}$ holds:
$\tfrac{d}{dt}H(x)$ $=\mathrm{grad}\,H^{\top}\dot{x}=(\mathbf{E}^{\star}\mathbf{z})^{\top}\dot{x}=\mathbf{z}^{\top}(\mathbf{J}-\mathbf{R})\mathbf{z}+\mathbf{z}^{\top}\mathbf{B}\mathbf{u}\leq\mathbf{y}^{\top}\mathbf{u}.$
The full set of Maxwell’s equations formulated with electrodynamic potentials already has a port-Hamiltonian structure (3) with $\mathbf{z}=\mathbf{x}:=(\tfrac{\partial}{\partial t}\bm{A},\varphi,\bm{H}\!)^{\top}$,
$\mathbf{E}_{\mathrm{Maxwell}}=\left[\begin{array}[]{ccc}\varepsilon&\varepsilon\mathrm{grad}&0\\ \mathrm{div}\varepsilon&\mathrm{div}\varepsilon\mathrm{grad}&0\\ 0&0&\mu\end{array}\right],$ $\mathbf{J}=\left[\begin{array}[]{ccc}0&0&-\mathrm{curl}\\ 0&0&0\\ \mathrm{curl}&0&0\end{array}\right],$ (5)
$\mathbf{R}=\left[\begin{array}[]{ccc}\kappa&\kappa\mathrm{grad}&0\\ \mathrm{div}\kappa&\mathrm{div}\kappa\mathrm{grad}&0\\ 0&0&0\end{array}\right],$ $\mathbf{B}=\left[\begin{array}[]{c}\!\mathbf{I}\!\!\\ \!\mathrm{div}\!\!\\ \!0\!\!\end{array}\right],$
where $\kappa,\varepsilon,\mu$ denote the coefficients for the electric conductivity, the permittivity and the permeability, respectively, $\bm{A},\varphi,\bm{H}$ the magnetic vector potential, the electric scalar potential and the magnetic field intensity. The input $\mathbf{u}(t)=\bm{j}_{\mathrm{s}}(t)$ is the source current densities and the output $\mathbf{y}(t)=\tfrac{\partial}{\partial t}\bm{A}+\mathrm{grad}{\varphi}(=-\bm{E})$ is the negative electric field intensity. The Hamiltonian of (5)
$H=\frac{1}{2}\left(\varepsilon\|\mathrm{grad}\varphi\|^{2}+\mu\|\bm{H}\|^{2}\right)+\varepsilon\frac{\partial}{\partial t}\bm{A}\cdot\mathrm{grad}\varphi+\varepsilon\|\frac{\partial}{\partial t}\bm{A}\|^{2},$ (6)
is the electromagnetic energy density and (4) holds with
$\mathrm{grad}\,H=\left(\varepsilon\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right),\,\mathrm{div}\left(\varepsilon\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)\right),\,\mu\bm{H}\right)^{\top}=\mathbf{E}_{\mathrm{Maxwell}}^{\star}{\mathbf{z}}.$
The resulting PHS inequality $\frac{d}{dt}H\leq{\mathbf{\bm{y}}}^{\top}{\mathbf{\bm{u}}}=\mathbf{j}_{\mathrm{s}}^{\top}(d_{t}\mathbf{A}+\mathrm{grad}{\varphi})$ corresponds to a discrete variant of Poynting’s theorem formulated without grid boundary radiation losses.
@@ 2 EMQS PHS Formulations
$\begin{split}\underbrace{\left[\begin{array}[]{ccc}0&\varepsilon\mathrm{grad}&0\\ 0&\mathrm{div}\varepsilon\mathrm{grad}&0\\ 0&0&\mu\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS1}}}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]&=\left[\left[\begin{array}[]{ccc}0&0&-\mathrm{curl}\\ 0&0&0\\ \mathrm{curl}&0&0\end{array}\right]\right.\\ &\phantom{abcd}-\left.\left[\begin{array}[]{ccc}\kappa&\kappa\mathrm{grad}&0\\ \mathrm{div}\kappa&\mathrm{div}\kappa\mathrm{grad}&0\\ 0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\!\\ \!\mathrm{div}\!\!\\ \!0\!\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (7)
with input $\mathrm{u}(t)=\bm{j}_{\mathrm{s}}$ and output $\mathrm{y}(t)=\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)=-\bm{E}.$ Equation (7) refers to a formulation without additional gauge equation yet, i.e., a Coulomb-type gauge condition
@@
The compatibility condition (4) can not hold as the first row of the conjugate operator matrix $\mathbf{E}_{\mathrm{EMQS1}}^{\star}$ with only zero entries would require that a Hamiltonian $H$ must not contain an expression $\frac{\partial}{\partial t}\bm{A}$, whereas the second row requires $\mathrm{div}(\varepsilon\frac{\partial}{\partial t}\bm{A})$ to occur in the Hamiltonian $H$. This argument also holds if a Coulomb-type gauge condition
$\mathrm{div}\left(\hat{\kappa}\frac{\partial}{\partial t}\bm{A}\right)=0$ (9)
is added to (2) as in [10] with an artificial electrical conductivity $\hat{\kappa}$ defined either in the physically non-conductive regions with $\kappa=0$ or throughout the computational domain, alternatively.
@@
$\begin{split}\underbrace{\left[\begin{array}[]{ccc}0&\varepsilon\mathrm{grad}&0\\ \mathrm{div}\varepsilon&\mathrm{div}\varepsilon\mathrm{grad}&0\\ 0&0&\mu\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS2}}}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]&=\left[\left[\begin{array}[]{ccc}0&0&-\mathrm{curl}\\ 0&0&0\\ \mathrm{curl}&0&0\end{array}\right]\right.\\ &-\left.\left[\begin{array}[]{ccc}\kappa&\kappa\mathrm{grad}&0\\ \mathrm{div}\kappa&\mathrm{div}\kappa\mathrm{grad}&0\\ 0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\!\\ \!\mathrm{div}\!\!\\ \!0\!\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (10)
with input $\mathbf{u}(t)=\mathbf{j}_{\mathrm{s}}$ and output $\mathbf{y}(t)=\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)=-\bm{E}.$ This approach implicitly enforces the condition $\mathrm{div}\left(\varepsilon\frac{\partial^{2}}{\partial t^{2}}\bm{A}\right)=0,$ which is the difference between the Darwin continuity equation (2) implicitly given via the Darwin-Ampère equation (1) and the full Maxwell continuity equation. Discrete versions of this symmetrized EMQS formulation are introduced in [4, 9].
The symmetrized EMQS formulation (10), however, yields a PHS formulation as the prerequisite compatibility relation (4)
$\begin{split}{\mathbf{E}}_{\mathrm{EMQS2}}^{\star}{\mathbf{z}}&=\left(\varepsilon\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right),\mathrm{div}\left(\varepsilon\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)\right),\mu\bm{H}\right)^{\top}\\ &=\mathrm{grad}\,H_{\mathrm{EMQS}}\end{split}$ (11)
@@
$H_{\mathrm{EMQS}}=\frac{1}{2}\left(\varepsilon\|\mathrm{grad}\varphi\|^{2}+\mu\|\bm{H}\|^{2}\right)+\varepsilon\frac{\partial}{\partial t}\bm{A}\cdot\mathrm{grad}\varphi,$ (12)
corresponding to the energy density of the EMQS field approximation.
A non-symmetric gauging, i.e., adding a non-physical Coulomb-type gauge expression $\mathrm{div}\left(\hat{\varepsilon}{\it d_{t}}^{2}\mathbf{a}\right)=0$ with a non-physical permittivity value $\hat{\varepsilon}$ to the the Darwin continuity equation (2) (similar to the introducing a non-physical conductivity $\kappa$ in the non-conductive regions in [10]) again results in a non-symmetric matrix $\mathbf{E}_{\mathrm{EMQS3}}$ which poses contradicting demands on a Hamiltonian expression if (4) were to hold, i.e., the corresponding EMQS formulation is not a PHS.

### v2

@@ 1 Introduction
Electromagnetic field problems, where objects are small in comparison to the shortest wave length of the problem, belong to the regime of quasistatic field models, where electromagnetic wave propagation effects are neglected. In addition to the established magneto- and electro-quasistatic field models, electromagneto-quasistatic field models, or so called Darwin-type field models w.r.t. to the original work in [4], have become of interest as quasistatic approximations of the full set of Maxwell equations.
These field models are of relevance for applications such as e.g. high-frequency coils or middle frequency transformers (MFTs), where resistive, capacitive and inductive field effects are considered simultaneously. The EMQS models of Darwin-type commonly feature the Darwin-Ampère equation formulated in terms of the magnetic vector potential $\bm{A}$ and the electric scalar potential $\varphi$ as
$\mathrm{curl}(\nu\mathrm{curl}\bm{A})+\kappa\frac{\partial}{\partial t}\bm{A}+\kappa\mathrm{grad}\varphi+\varepsilon\mathrm{grad}\frac{\partial}{\partial t}\varphi=\bm{j}_{\mathrm{s}},$ (1)
where $\kappa,\varepsilon,\nu$ denote the coefficients for the electric conductivity, the permittivity and the reluctivity, respectively.
@@
Other EQMS formulations combine (1) with the Maxwell continuity equation [11, 8] which yields symmetric formulations. Combined with the finite integration technique (FIT) [16] or related mimetic discretization schemes this formulation implicitly enforces a Coulomb-type gauge $\mathrm{div}(\varepsilon\frac{\partial}{\partial t}\bm{A})=0.$ Related EMQS formulations as in [17] additionally and explicitly enforce this Coulomb-type gauge within Lagrange multiplier formulations. A number of EMQS formulations use domain dependent gauge equations and split scalar potentials to enforce a low-frequency stabilization [1, 11].
Port-Hamiltonian systems (PHS) models [10] are used to maintain physical properties such as energy conservation and/or dissipative inequalities. While this contribution focuses on PHS formulations for EMQS field models, there is also interest in PHS in other areas of electromagnetic field research, such as PHS for Maxwell equations, as outlined in references [5] and [6], and PHS for magneto-quasistatics, as presented in references [14] and [15].
We consider port-Hamiltonian partial differential-algebraic equation (pH-PDAE) system of the form [2]
${\mathbf{E}}\,\frac{d}{dt}\mathbf{x}=({\mathbf{J}}-{\mathbf{R}})\,{\mathbf{z}}(\mathbf{x})+{\mathbf{B}}\,{\mathbf{u}}(t),\quad\mathbf{x}(0)=\mathbf{x}_{0},\quad{\mathbf{y}}(t)={\mathbf{B}}^{\star}\mathbf{z}(\mathbf{x}).$ (3)
[…]
$\langle Ex,y\rangle$ $=\langle x,E^{\star}y\rangle,\quad\langle Rx,y\rangle=\langle x,R^{\star}y\rangle,\quad\langle Rx,x\rangle\geq 0.$
In addition, we assume that the differential operator ${\mathbf{J}}$ is formally skew-adjoint.
The function $\mathbf{z}$ depends on the inner state variable $\mathbf{x}$ and the input is given by $\mathbf{u}(t)$. If the compatibility relation
@@
holds for the Hamiltonian $H=H(\mathbf{x})$, which represents an energy expression, the dissipativity inequality $\frac{d}{dt}H(\mathbf{x})\leq{\mathbf{y}}^{\top}{\mathbf{u}}$ holds:
$\tfrac{d}{dt}H(\mathbf{x})$ $=\color[rgb]{0.0,0.0,0.0}\langle\mathrm{grad}\,H,\dot{\mathbf{x}}\rangle=\langle\mathbf{E}^{\star}\mathbf{z},\dot{\mathbf{x}}\rangle=\underbrace{\langle\mathbf{z},\mathbf{J}\mathbf{z}\rangle}_{=0}-\underbrace{\langle\mathbf{z},\mathbf{R}\mathbf{z}\rangle}_{\geq 0}+\langle\underbrace{\mathbf{B}^{\star}\mathbf{z}}_{\mathbf{y}},\mathbf{u}\rangle\leq\langle\mathbf{y},\mathbf{u}\rangle,$ (5)
with $\mathrm{grad}\,H$ given as the variational derivative of the Hamiltonian [14, Sect. 2.3]. The full set of Maxwell’s equations formulated with electrodynamic potentials already has a port-Hamiltonian structure (3) with $\mathbf{z}=\mathbf{x}:=(\tfrac{\partial}{\partial t}\bm{A},\varphi,\bm{H}\!)^{\top}$,
$\mathbf{E}_{\mathrm{Maxwell}}=\left[\begin{array}[]{ccc}\varepsilon&\varepsilon\mathrm{grad}&0\\ -\mathrm{div}\varepsilon&-\mathrm{div}\varepsilon\mathrm{grad}&0\\ 0&0&\mu\end{array}\right],$ $\mathbf{J}=\left[\begin{array}[]{ccc}0&0&-\mathrm{curl}\\ 0&0&0\\ \mathrm{curl}&0&0\end{array}\right],$ (6)
$\mathbf{R}=\left[\begin{array}[]{ccc}\kappa&\kappa\mathrm{grad}&0\\ -\mathrm{div}\kappa&-\mathrm{div}\kappa\mathrm{grad}&0\\ 0&0&0\end{array}\right],$ $\mathbf{B}=\left[\begin{array}[]{c}\!1\!\!\\ \!-\mathrm{div}\!\!\\ \!0\!\!\end{array}\right],$
with suitably chosen boundary condtions, and where $\mu=\nu^{-1}$ denotes the coefficient for the magnetic permeability. $\bm{A},\varphi,\bm{H}$ denote the magnetic vector potential, the electric scalar potential and the magnetic field intensity, respectively. The input $\mathbf{u}(t)=\bm{j}_{\mathrm{s}}(t)$ is the source current densities and the output $\mathbf{y}(t)=\tfrac{\partial}{\partial t}\bm{A}{\color[rgb]{0.0,0.0,0.0}+\mathrm{grad}\varphi}(=-\bm{E})$ is the negative electric field intensity. The initial value $\mathbf{x}_{0}:=(\tfrac{\partial}{\partial t}\bm{A}_{0},\varphi_{0},\bm{H}_{0}\!)^{\top}$ must satisfy the constraint $\mu\mathbf{H}_{0}=\mathrm{curl}\mathbf{A}_{0}$. Furthermore, if time discretization is employed, an analogous constraint must hold in a time discretized sense [3].
The Hamiltonian of (6)
[…]
$\frac{d}{dt}H$ $=\langle\varepsilon(\mathrm{grad}\varphi+\bm{\dot{A}}),\mathrm{grad}\dot{\varphi}\rangle+\langle\mu\bm{H},\bm{\dot{H}}\rangle+\langle\varepsilon\bm{\ddot{A}},\bm{\dot{A}}+\mathrm{grad}\varphi\rangle$ (8)
$=\langle\mathrm{grad}H,\bm{\dot{x}}\rangle$ (9)
with
$\mathrm{grad}H$ $=\begin{bmatrix}\varepsilon(\mathrm{grad}\varphi+\bm{\dot{A}})\\ -\mathrm{div}\varepsilon(\mathrm{grad}\varphi+\bm{\dot{A}})\\ \mu\bm{H}\end{bmatrix}.$ (10)
One easily verfies the compatability condition (4), as we get $\mathbf{E}_{\mathrm{Maxwell}}^{\star}{\mathbf{z}}=\mathbf{E}_{\mathrm{Maxwell}}{\mathbf{z}}=\mathrm{grad}H.$ The resulting PHS inequality $\frac{d}{dt}H\leq\color[rgb]{0.0,0.0,0.0}\langle\mathbf{\bm{y}},{\mathbf{\bm{u}}}\rangle=-\langle\bm{E},\mathbf{j}_{\mathrm{s}}\rangle$ corresponds to a variant of Poynting’s theorem formulated without boundary radiation losses $\mathrm{div}\mathbf{S}=\mathrm{div}(\bm{E}\times\bm{H})$.
@@ 2 EMQS PHS Formulations
$\begin{split}\underbrace{\left[\begin{array}[]{ccc}0&\varepsilon\mathrm{grad}&0\\ 0&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon\mathrm{grad}&0\\ 0&0&\mu\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS1}}}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]&=\left[\left[\begin{array}[]{ccc}0&0&-\mathrm{curl}\\ 0&0&0\\ \mathrm{curl}&0&0\end{array}\right]\right.\\ &\phantom{abcd}-\left.\left[\begin{array}[]{ccc}\kappa&\kappa\mathrm{grad}&0\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa\mathrm{grad}&0\\ 0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\!\\ \!\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\!\!\\ \!0\!\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (11)
with input $\mathbf{u}(t)=\bm{j}_{\mathrm{s}}$ and output $\mathbf{y}(t)=\left(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi\right)=-\bm{E}.$ Equation (11) refers to a formulation without additional gauge equation yet, i.e., a Coulomb-type gauge condition
@@
In this case, $\mathbf{E}_{\mathrm{EMQS1}}$ is not self-adjoint. In addition, the compatibility condition (4) can not hold as the first row of the conjugate operator matrix $\mathbf{E}_{\mathrm{EMQS1}}^{\star}$ with only zero entries would require that a Hamiltonian $H$ must not contain an expression $\frac{\partial}{\partial t}\bm{A}$, whereas the second row requires $\mathrm{div}(\varepsilon\frac{\partial}{\partial t}\bm{A})$ to occur in the Hamiltonian $H$. This argument also holds if a Coulomb-type gauge condition $\mathrm{div}\left(\hat{\kappa}\frac{\partial}{\partial t}\bm{A}\right)=0$ is added to (2) as in [13] with an artificial electrical conductivity $\hat{\kappa}$ defined either in the physically non-conductive regions with $\kappa=0$ or throughout the computational domain, alternatively.
@@
$\begin{split}\underbrace{\left[\begin{array}[]{ccc}0&\varepsilon\mathrm{grad}&0\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\varepsilon\mathrm{grad}&0\\ 0&0&\mu\end{array}\right]}_{{\mathbf{E}}_{\mathrm{EMQS2}}}\frac{\partial}{\partial t}\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]&=\left[\left[\begin{array}[]{ccc}0&0&-\mathrm{curl}\\ 0&0&0\\ \mathrm{curl}&0&0\end{array}\right]\right.\\ &-\left.\left[\begin{array}[]{ccc}\kappa&\kappa\mathrm{grad}&0\\ \color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa&\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\kappa\mathrm{grad}&0\\ 0&0&0\end{array}\right]\right]\left[\begin{array}[]{c}\!\frac{\partial}{\partial t}\bm{A}\!\\ \!\varphi\!\\ \!\bm{H}\!\end{array}\right]\!\!+\!\!\left[\begin{array}[]{c}\!1\!\!\\ \!\color[rgb]{0.0,0.0,0.0}-\mathrm{div}\!\!\\ \!0\!\!\end{array}\right]\bm{j}_{\mathrm{s}},\end{split}$ (13)
with input $\mathbf{u}(t)=\mathbf{j}_{\mathrm{s}}$ and output $\mathbf{y}(t)=(\frac{\partial}{\partial t}\bm{A}+\mathrm{grad}\varphi)=-\bm{E}.$ This approach implicitly enforces the condition $\mathrm{div}(\varepsilon\frac{\partial^{2}}{\partial t^{2}}\bm{A})=0,$ which is the difference between the Darwin continuity equation (2) implicitly given via the Darwin-Ampère equation (1) and the full Maxwell continuity equation. Discrete versions of this symmetrized EMQS formulation are introduced in [8, 11].
The symmetrized EMQS formulation (13), however, yields a PHS formulation as the prerequisite compatibility relation (4) ${\mathbf{E}}_{\mathrm{EMQS2}}^{\star}{\mathbf{z}}=\mathrm{grad}\,H_{\mathrm{EMQS}}$ with
$\mathrm{grad}H_{\mathrm{EMQS}}$ $=\begin{bmatrix}\varepsilon\mathrm{grad}\varphi\\ -\mathrm{div}\varepsilon\frac{\partial}{\partial t}\bm{A}\color[rgb]{0,0,0}-\mathrm{div}\varepsilon\mathrm{grad}\varphi\\ \mu\bm{H}\end{bmatrix}$
@@
$H_{\mathrm{EMQS}}=\int_{\Omega}\left\{\color[rgb]{0,0,0}\frac{1}{2}\left(\varepsilon\|\mathrm{grad}\varphi\|^{2}+\mu\|\bm{H}\|^{2}\right)+\varepsilon\frac{\partial}{\partial t}\bm{A}\cdot\mathrm{grad}\varphi\right\},$ (14)
corresponding to the energy of the EMQS field approximation.
A non-symmetric gauging, i.e., adding a non-physical Coulomb-type gauge expression $\mathrm{div}\left(\hat{\varepsilon}\frac{\partial^{2}}{\partial t^{2}}\mathbf{a}\right)=0$ with a non-physical permittivity value $\hat{\varepsilon}$ to the Darwin continuity equation (2) (similar to the introducing a non-physical conductivity $\kappa$ in the non-conductive regions in [13]) again results in a non-self-adjoint matrix $\mathbf{E}_{\mathrm{EMQS3}}$ which poses contradicting demands on a Hamiltonian expression if, given $\mathbf{z}=\mathbf{x}$, (4) were to hold.